# Charactarisation of distal actions of automorphisms on the space of one-parameter subgroups of Lie groups

### v1

@@ 3. Characterisation of distal actions on the space of one-parameter subgroups
The notation is inspired by the fact that the class ${\mathcal{C}^{\prime}}$ is a subset of the class $\mathcal{C}$ defined by Dani and McCrudden in [13]. The latter consists of connected Lie groups which admit a (real) linear representation with discrete kernel. Proposition 2.5 of [13] shows that $G$ is in class $\mathcal{C}$ if and only if for the radical $R$ of $G$, the intersection of $\overline{[R,R]}$ with the center of $G$ has no non-trivial compact subgroup. Therefore class ${\mathcal{C}^{\prime}}$ is a subset of class $\mathcal{C}$. Note that any connected Lie group $G$ admits a maximal compact connected central subgroup $M$, and $G/M$ has no non-trivial compact central subgroup and it belongs to class ${\mathcal{C}^{\prime}}$.
Recall that for a connected Lie group $G$ with the Lie algebra ${\mathfrak{g}}$, there is an exponential map $\exp:{\mathfrak{g}}\to G$, and for any $T\in{\rm Aut}(G)$, there is a Lie algebra automorphism ${\rm d\,}T:{\mathfrak{g}}\to{\mathfrak{g}}$ such that $\exp(({\rm d\,}T)(X))=T(\exp X)$, $X\in{\mathfrak{g}}$, and $T\mapsto{\rm d\,}T$ is a homomorphism and homeomorphism from ${\rm Aut}(G)$ onto its image in ${\rm GL}({\mathfrak{g}})$ which is closed; the image consists of Lie algebra automorphisms of ${\mathfrak{g}}$. If $G$ is in class ${\mathcal{C}^{\prime}}$, then ${\rm Aut}(G)$ is almost algebraic as a subgroup of ${\rm GL}({\mathfrak{g}})$, it is closed and has finitely many connected components (cf. [12]). If $H$ is a closed connected normal subgroup of $G$ with the Lie algebra ${\mathfrak{h}}$ which is a Lie subalgebra of ${\mathfrak{g}}$ and it is an ideal in ${\mathfrak{g}}$. Moreover, the Lie algebra of $G/H$ is isomorphic to ${\mathfrak{g}}/{\mathfrak{h}}$. Also, if $H$ is $T$-invariant and if $\overline{T}$ denotes the automorphism of $G/H$ induced by $T$. Then ${\rm d\,}\overline{T}=\overline{{\rm d\,}T}$, which is the Lie algebra automorphism on ${\mathfrak{g}}/{\mathfrak{h}}$ induced by ${\rm d\,}T$. Also ${\rm d\,}T|_{H}$ is the same as $({\rm d\,}T)|_{\mathfrak{h}}$. Note that the eigenvalues of ${\rm d\,}\overline{T}$ (as well as those of ${\rm d\,}T|_{H}$) are eigenvalues of ${\rm d\,}T$.
It is known that $T$ acts distally on $G$ if and only if ${\rm d\,}T$ acts distally on ${\mathfrak{g}}$, and the latter statement is equivalent to the statement that all the eigenvalues of ${\rm d\,}T$ have absolute value 1 (cf [2]). It is also known that if $G$ admits an automorphism $T$ such that ${\rm d\,}T$ has no eigenvalue of absolute value 1 (equivalently, if $T$ is expansive; see [37], see also [7]), then $G$ must be nilpotent (this is well-known, see e.g. [37]). We will need the following lemma which is slightly stronger. The lemma may be known, but we give a short proof for the sake of completeness.
@@
Let $M$ be the largest compact connected central subgroup of $G$. Then $G/M$ is in class ${\mathcal{C}^{\prime}}$. As $G$ is not nilpotent, $G/M$ is also not nilpotent. Since $M$ is characteristic in $G$, we may replace $G$ by $G/M$ and assume that $G\in{\mathcal{C}^{\prime}}$, i.e. it has no compact central subgroup of positive dimension. If $T$ has finite order, then $({\rm d\,}T)^{n}={\rm Id}$ for some $n\in{\mathbb{N}}$, and the assertion follows in this case. Now suppose $T$ does not have finite order.
@@
Note that the largest compact central subgroup and its connected component of the identity in a connected Lie group are characteristic. We note the following useful proposition which will enable us to work on Lie groups in class ${\mathcal{C}^{\prime}}$.
Proposition 3.4.
Proof.
We also need the following elementary lemma.
Lemma 3.5.
Let $G$ be a simply connected nilpotent Lie group with the Lie algebra ${\mathfrak{g}}$ and let $T\in{\rm Aut}(G)$. Then the set ${\mathcal{P}}_{1}(G)$ consisting of all closed one-parameter subgroups of $G$ is closed in ${\rm Sub}_{G}$ and is the same as ${\rm Sub}^{p}_{G}$. Moreover, $T$ acts distally on ${\rm Sub}^{p}_{G}$ if and only if $dT$ acts distally on ${\rm Sub}^{p}_{\mathfrak{g}}$.
Proof.
In a simply connected nilpotent group, all one-parameter subgroups are closed and there is a one-to-one correspondence between non-trivial (closed) one-parameter subgroups of $G$ and one-dimensional subspaces of ${\mathfrak{g}}$. The correspondence is via the exponential map $\exp:{\mathfrak{g}}\to G$ with $\log$ as its inverse. Namely, for $\{x_{t}\}$, $x_{t}=\exp tX$, for some $X\in{\mathfrak{g}}$, and $\log x_{t}=tX$, $t\in{\mathbb{R}}$. Suppose $\{H_{n}\}\subset{\mathcal{P}}_{1}(G)$ is such that $H_{n}\to H$ in ${\rm Sub}_{G}$. Then $\log H_{n}\in{\mathcal{P}}_{1}({\mathfrak{g}})$ and using Lemma 2.1 and the continuity of the exponential map and its inverse, it is easy to show that $\{\log H_{n}\}$ converges to (say) $B$ in ${\rm Sub}^{p}_{\mathfrak{g}}$. and since ${\mathfrak{g}}$ is isomorphic to ${\mathbb{R}}^{n}$ for some $n$, ${\rm Sub}^{p}_{\mathfrak{g}}={\mathcal{P}}_{1}({\mathfrak{g}})$ which is closed and we have that $B\in{\mathcal{P}}_{1}({\mathfrak{g}})$. Then $H=\exp B\in{\mathcal{P}}_{1}(G)$ and, in particular, we get that ${\mathcal{P}}_{1}(G)$ is closed in ${\rm Sub}_{G}$, i.e. ${\mathcal{P}}_{1}(G)={\rm Sub}^{p}_{G}$. In particular, $H_{n}\to H$ in ${\rm Sub}^{p}_{G}$ implies that $\log H_{n}\to\log H$ in ${\rm Sub}^{p}_{\mathfrak{g}}$.
Proof of Theorem 1.1 $(1)$ and $(2)$ for nilpotent groups.
Let $G$ be a (non-trivial) connected Lie group in class ${\mathcal{C}^{\prime}}$, i.e. $G$ does not have any non-trivial compact connected central subgroup. Let $T\in{\rm Aut}(G)$.
[…]
As $G$ is simply connected and nilpotent, by Lemma 3.5, ${\rm d\,}T$ acts distally on ${\rm Sub}^{p}_{\mathfrak{g}}$. Then from (1) above, we have that ${\rm d\,}T\in{\mathcal{K}}\{r{\rm Id}\mid r\neq 0\}$ for some compact subgroup ${\mathcal{K}}$ of ${\rm GL}({\mathfrak{g}})$ and ${\rm Id}$ is the identity map on ${\mathfrak{g}}$. As ${\rm d\,}T^{n}=\psi\,\circ\,r\,{\rm Id}$ for some $r\in{\mathbb{R}}\setminus\{0\}$ and $\psi\in{\mathcal{K}}$. If $r=\pm 1$, then ${\rm d\,}T^{2}\in{\mathcal{K}}$ and $T$ generates a relatively compact subgroup of ${\rm Aut}(G)$. If possible suppose $r\neq\pm 1$. Since ${\rm Aut}(G)$ is almost algebraic, there exists $n\in{\mathbb{N}}$ such that ${\rm d\,}T^{n}$ is contained in a connected almost algebraic subgroup ×${\mathbb{R}}\times\mathbb{T}^{k}$ of ${\rm Aut}(G)^{0}$ (which is identified with its image in ${\rm GL}({\mathfrak{g}})$), where $\mathbb{T}^{k}=(\mathbb{S}^{1})^{k}$ for some $k\in{\mathbb{N}}$. Therefore, $\psi^{n}\,\circ\,r^{n}\,{\rm Id}\in{\mathbb{R}}\times\mathbb{T}^{k}\subset{\rm Aut}(G)^{0}$. Hence both $\psi^{n}$ and $r^{n}\,{\rm Id}\in{\rm Aut}(G)^{0}$. But $r^{n}\,{\rm Id}$ is not a Lie algebra automorphism of ${\mathfrak{g}}$ as $r^{n}\neq 1$, for if $r^{n}[X,Y]=r^{n}{\rm Id}[X,Y]=[r^{n}{\rm Id}(X),r^{n}{\rm Id}(Y)]=r^{2n}[X,Y]$, then $[X,Y]=0$ for all $X,Y\in{\mathfrak{g}}$, which leads to a contradiction as $G$ is not abelian. Thus, $r=\pm 1$ and $T$ is contained in a compact subgroup of ${\rm Aut}(G)$. The converse statement is obvious. Thus (2) holds if $G$ is nilpotent. ∎
We will continue with the proof of Theorem 1.1 after proving Theorems 1.2 and 1.3 and some elementary results.
For a Lie group $G$ and $T\in{\rm Aut}(G)$, let $C(T)$ denote the contraction group of $T$ and it is defined as
@@
Note that $C(T)$ is a connected nilpotent subgroup of $G$ and it is simply connected if it is closed in $G$ (see [41], see also [20]). Both $C(T)$ and $C(T^{-1})$ are trivial if and only if $T$ acts distally on $G$ (cf.  [34], Theorem 4.1). For a connected Lie group $G$, $C(T)$ is trivial if and only if all the eigenvalues of ${\rm d\,}T$ have absolute value greater than or equal to 1 ([34, 2]).
We need the following lemma which can be proven easily by using Theorem 1.1 of [14], Theorem 2.4 and Corollary 2.7 of [20]. Note that the largest compact connected central subgroup of $G$ is characteristic in $G$. The lemma generalises a part of the statement in Proposition 4.3 of [34].
@@
Let $M$ be the largest compact connected central subgroup of $G$. Let $\pi:G\to G/M$ be the natural projection and let $T^{\prime}\in{\rm Aut}(G/M)$ be the automorphism induced by $T$ on $G/M$. As $G/M$ does not have any non-trivial compact connected central subgroup, using Theorem 1.1 of [14], one can show that $C(T^{\prime})$ is closed (see the proof of Proposition 4.3 of [34] for more details). Hence $\pi^{-1}(C(T^{\prime}))=C_{M}(T):=\{x\in G\mid T^{n}(x)M\to M\mbox{ in }G/M\}$ is closed. Suppose (1) holds, i.e. $T$ acts distally on $M$. Then $C(T)\cap M=C(T|_{M})=\{e\}$. As $C_{M}(T)$ is closed, by Corollary 2.7 of [20], $C(T)$ is closed. Similarly, $C(T^{-1})$ is closed. The rest of the assertion in (2) follows from Corollary 2.4 in [41]. This proves $(1)\implies(2)$.
@@
Let $X\in{\mathfrak{g}}$, $X\neq 0$ be such that ${\rm d\,}T(X)=X$ and let $A=\{\exp tX\}$, a one-parameter subgroup, it is either closed or its closure $\bar{A}$ is compact. Since $T|_{A}={\rm Id}$, we have that $T|_{\bar{A}}={\rm Id}$ and it follows that $\bar{A}$ normalises $C(T)$ and $\bar{A}\cap C(T)=\{e\}$. Let $H=\overline{A\ltimes C(T)}$. Then $H$ is solvable. As $A$ normalises $C(T)$, so does $H$. Now $H/C(T)$ is connected and abelian as it is the closure of the image of the one-parameter subgroup $A$ in it. Then $H/C(T)$ is either isomorphic to ${\mathbb{R}}$ or it is compact. In the first case it follows that $H=A\ltimes C(T)$. In the second case, since $C(T)$ is a vector group, $H=K\ltimes C(T)$ for some (maximal) compact subgroup $K\subset H$ (cf. [24], Lemma 3.7). Here, $H$ is a closed $T$-invariant subgroup of $G$ and $T(h)\in h\,C(T)$ for all $h\in H$. If $A$ (resp. $\bar{A}$) is compact, then $H=A\ltimes C(T)$ (resp. $H=\bar{A}\ltimes C(T)$). Note that $C(T)$ is isomorphic to ${\mathbb{R}}^{d}$ and $T_{1}$, the restriction of $T$ to $C(T)$, is as described earlier.
Step 3: Suppose $H/C(T)$ is isomorphic to ${\mathbb{R}}$. Then so is $A$ and $H=A\ltimes C(T)$. Let $U,V$ be open neighbourhoods of the identity $e$ in $H$ such that $V^{2}\subset U$ and every element of $U$ is contained in a one-parameter subgroup. Let $x=ac$ for some $a\in A\cap V$ and $c\in C(T)\cap V$ with $c\neq e$. Then $x\in V^{2}\subset U$ and $T(x)=xy$ for $y=c^{-1}T(c)\in C(T)$ and $y\neq e$ as $c\in C(T)$ and $c\neq e$. Now $x$ is contained in a one-parameter subgroup (say) $\{x_{t}\}$ in $H$ as $x=x_{1}$. Then $T(x_{t})=x_{t}y_{t}$, $y_{t}\in C(T)$, $t\in{\mathbb{R}}$ and $y_{1}=y\neq e$. Here, $\{x_{t}\}$ is closed as $H$ is simply connected and solvable. As ${\rm Sub}^{p}_{H}$ is compact, we get that $T^{n_{k}}(\{x_{t}\})\to B$ for some $B\in{\rm Sub}^{p}_{H}$ and some unbounded sequence $\{n_{k}\}\subset{\mathbb{N}}$. By Lemma 2.2, $B\neq\{e\}$. Let $b\in B\setminus\{e\}$. Then $T^{n_{k}}(x_{t_{k}})\to b$ for some sequence $\{t_{k}\}\subset{\mathbb{R}}$. Here, $\{x_{t}\}C(T)/C(T)=H/C(T)$ is isomorphic to ${\mathbb{R}}$. Since $x_{t_{k}}C(T)\to b\,C(T)$ in $H/C(T)$, we get that $\{t_{k}\}$ is bounded. Passing to a subsequence if necessary, we may assume that $t_{k}\to t_{0}$. Then $x_{t_{k}}\to x_{t_{0}}$ and $y_{t_{k}}\to y_{t_{0}}$. As all the eigenvalues of $T|_{C(T)}=T_{1}=\phi\circ\alpha\,I$ have absolute value $\alpha$ with $10<\alpha<$ 1, we get that $I-T_{1}$ is invertible in ${\rm GL}(d,{\mathbb{R}})$, where $I$ is the identity matrix in ${\rm GL}(d,{\mathbb{R}})$. Using the addition as the group operation in ${\mathbb{R}}^{d}$, and the fact that $T_{1}^{n}(y)\to 0$ for all $y\in{\mathbb{R}}^{d}$, we get that
@@
Step 4: Now we have that $H/C(T)$ is compact and $H=K\ltimes C(T)$, where $K$ is a maximal compact subgroup of $H$. We show that $T(x)=x$ for all $x\in K$. We know that $T(x)\in x\,C(T)$ for all $x\in K$. Since $K$ is connected and abelian, the set $F$ of roots of unity in $K$ is dense in $K$ and it is enough to show that $T(x)=x$ for all $x\in F$. If possible, suppose there exists $x\in F$ such that $T(x)\neq x$. Let $m\in{\mathbb{N}}$ be the order of $x$, i.e. $x^{m}=e$ and $m\neq 1$. Let $\{x_{t}\}$ be a one-parameter subgroup in $K$ such that $x_{1}=x$. As $x_{1}^{m}=e$, $\{x_{t}\}$ is closed and isomorphic to ${\mathbb{S}}^{1}$ and $T(x_{t})=x_{t}y_{t}$, where $y_{t}\in C(T)$, $t\in{\mathbb{R}}$ and $y_{1}=x^{-1}T(x)\neq e$. There exists an unbounded sequence $\{n_{k}\}\subset{\mathbb{N}}$ such that $T^{n_{k}}(\{x_{t}\})\to B$ for some $B\subset H$. Then $B\neq\{e\}$ as $x_{1}\neq e$ and $\{x_{t}\}$ is connected (cf. Lemma 2.2). Let $b\in B$ be such that $b\neq e$. Then $T^{n_{k}}(x_{t_{k}})\to b$ for some $\{t_{k}\}\subset{\mathbb{R}}$. Here, $\{x_{t}\}$ is periodic i.e. $x_{mn}=e$ and $x_{m+t}=x_{t}$ for all $n\in{\mathbb{Z}}$ and $t\in{\mathbb{R}}$, where $m$ is the order of $x_{1}=x$. Hence we can choose $\{t_{k}\}$ such that it is bounded. Then $\{x_{t_{k}}\}$, and hence, $\{y_{t_{k}}\}$ is bounded. Now arguing as in Step 3, we get that $T(b)=b$. Therefore, $T(B)=B$, and hence $T^{n_{k}}(T(\{x_{t}\}))\to B$. Since $T$ acts distally on ${\rm Sub}^{p}_{G}$, we get that $T(\{x_{t}\})=\{x_{t}\}$. As $T$ is an automorphism and $\{x_{t}\}$ is isomorphic to ${\mathbb{S}}^{1}$, we get that $T(x_{t})=x_{t}$ or $T(x_{t})=x_{t}^{-1}=x_{-t}$. In either case, we get that $y_{1}\in K\cap C(T)=\{e\}$ This leads to a contradiction. Hence $T(x)=x$, for all $x\in F$, and hence for all $x\in K$.

### v2

@@ 3. Some properties of automorphisms and their actions on $G$ and ${\mathrm{Sub}}^{p}_{G}$
The notation is inspired by the fact that the class ${\mathcal{C}^{\prime}}$ is a subset of the class $\mathcal{C}$ defined by Dani and McCrudden in [13]. The latter consists of connected Lie groups which admit a (real) linear representation with discrete kernel. Proposition 2.5 of [13] shows that $G$ is in class $\mathcal{C}$ if and only if for the radical $R$ of $G$, the intersection of $\overline{[R,R]}$ with the center of $G$ has no non-trivial compact subgroup. Therefore class ${\mathcal{C}^{\prime}}$ is a subset of class $\mathcal{C}$. Note that any connected Lie group $G$ admits a maximal compact connected central subgroup $M$, and $G/M$ has no non-trivial compact central subgroup and hence it belongs to class ${\mathcal{C}^{\prime}}$. We may recall a result of Dani in [12] that if $G$ belongs to class ${\mathcal{C}^{\prime}}$, then ${\rm Aut}(G)$ is almost algebraic, i.e. it is an algebraic subgroup of ${\rm GL}({\mathfrak{g}})$, and in particular it has finitely many connected components.
It is known that $T\in{\rm Aut}(G)$ acts distally on $G$ if and only if ${\rm d\,}T$ acts distally on the Lie algebra ${\mathfrak{g}}$ of $G$, and the latter statement is equivalent to the statement that all the eigenvalues of ${\rm d\,}T$ have absolute value 1 (cf [2]). It is also known that if $G$ admits an automorphism $T$ such that ${\rm d\,}T$ has no eigenvalue of absolute value 1 (equivalently, if $T$ is expansive; see [38], and also [7]), then $G$ must be nilpotent (this is well-known, see e.g. [38]). We will need the following lemma which is slightly stronger. The lemma may be known, but we give a short proof for the sake of completeness.
@@
Let $M$ be the largest compact connected central subgroup of $G$. Then $G/M$ is in class ${\mathcal{C}^{\prime}}$. As $G$ is not nilpotent, $G/M$ is also not nilpotent. Since $M$ is characteristic in $G$, we may replace $G$ by $G/M$ and assume that $G\in{\mathcal{C}^{\prime}}$. If $T$ has finite order, then $({\rm d\,}T)^{n}={\rm Id}$ for some $n\in{\mathbb{N}}$, and the assertion follows in this case. Now suppose $T$ does not have finite order.
@@
For a connected Lie group $G$ and $T\in{\rm Aut}(G)$, let $C(T)$ denote the contraction group of $T$ and it is defined as
@@
Note that $C(T)$ is a connected nilpotent subgroup of $G$ and it is simply connected if it is closed in $G$ (see [43], see also [21]). Both $C(T)$ and $C(T^{-1})$ are trivial if and only if $T$ acts distally on $G$ (cf.  [35], Theorem 4.1). Note also that $C(T)$ is trivial if and only if all the eigenvalues of ${\rm d\,}T$ have absolute value greater than or equal to 1 ([35, 2]).
We will need the following lemma which can be proven easily by using Theorem 1.1 of [14], Theorem 2.4 and Corollary 2.7 of [21]; we will give a sketch of proof. Note that the largest compact connected central subgroup of $G$ is characteristic in $G$. The lemma generalises a part of the statement in Proposition 4.3 of [35].
@@
Let $M$ be the largest compact connected central subgroup of $G$. If $G=M$, then it is easy to see that $(1-3)$ are equivalent. Now suppose $G\neq M$. Let $\pi:G\to G/M$ be the natural projection and let $T^{\prime}\in{\rm Aut}(G/M)$ be the automorphism induced by $T$ on $G/M$. As $G/M\in{\mathcal{C}^{\prime}}$, using Theorem 1.1 of [14], one can show that $C(T^{\prime})$ is closed (see the proof of Proposition 4.3 of [35] for more details). Hence $\pi^{-1}(C(T^{\prime}))=C_{M}(T):=\{x\in G\mid T^{n}(x)M\to M\mbox{ in }G/M\}$ is closed. Suppose (1) holds, i.e. $T$ acts distally on $M$. Then $C(T)\cap M=C(T|_{M})=\{e\}$. As $C_{M}(T)$ is closed, by Corollary 2.7 of [21], $C(T)$ is closed. Similarly, $C(T^{-1})$ is closed. The rest of the assertion in (2) follows from Corollary 2.4 in [43]. This proves $(1)\implies(2)$.
@@
We note the following useful proposition which will enable us to work on the quotient Lie group which belongs to class ${\mathcal{C}^{\prime}}$.
Proposition 3.5.
Proof.
We will also need the following elementary lemma.
Lemma 3.6.
Let $G$ be a simply connected nilpotent Lie group with the Lie algebra ${\mathfrak{g}}$ and let $T\in{\rm Aut}(G)$. Then the set ${\mathcal{P}}_{1}(G)$ consisting of all closed one-parameter subgroups of $G$ is closed in ${\rm Sub}_{G}$ and is the same as ${\rm Sub}^{p}_{G}$. Moreover, $T$ acts distally on ${\rm Sub}^{p}_{G}$ if and only if ${\rm d\,}T$ acts distally on ${\rm Sub}^{p}_{\mathfrak{g}}$.
Proof.
In a simply connected nilpotent group, all one-parameter subgroups are closed and there is a one-to-one correspondence between non-trivial (closed) one-parameter subgroups of $G$ and one-dimensional subspaces of ${\mathfrak{g}}$. The correspondence is via the exponential map $\exp:{\mathfrak{g}}\to G$ with $\log$ as its inverse. Namely, for $\{x_{t}\}$, $x_{t}=\exp tX$, for some $X\in{\mathfrak{g}}$, and $\log x_{t}=tX$, $t\in{\mathbb{R}}$. Suppose $\{H_{n}\}\subset{\mathcal{P}}_{1}(G)$ is such that $H_{n}\to H$ in ${\rm Sub}_{G}$. Then $\log H_{n}\in{\mathcal{P}}_{1}({\mathfrak{g}})$ and using Lemma 2.1 and the continuity of the exponential map and its inverse, it is easy to show that $\{\log H_{n}\}$ converges to (say) $B$ in ${\rm Sub}^{p}_{\mathfrak{g}}$. Since ${\mathfrak{g}}$ is isomorphic to ${\mathbb{R}}^{n}$ for some $n$, ${\rm Sub}^{p}_{\mathfrak{g}}={\mathcal{P}}_{1}({\mathfrak{g}})$ which is closed and we have that $B\in{\mathcal{P}}_{1}({\mathfrak{g}})$. Then $H=\exp B\in{\mathcal{P}}_{1}(G)$, and we get that ${\mathcal{P}}_{1}(G)$ is closed in ${\rm Sub}_{G}$, i.e. ${\mathcal{P}}_{1}(G)={\rm Sub}^{p}_{G}$. Therefore, $H_{n}\to H$ in ${\rm Sub}^{p}_{G}$ implies that $\log H_{n}\to\log H$ in ${\rm Sub}^{p}_{\mathfrak{g}}$.
4. Proofs of the main results
In this section, we prove the main results stated in the introduction. Additionally, we get some corollaries which are derived from the main results. We give some examples of Lie groups $G$ with a nontrivial central torus on which all automorphisms act distally, which is stronger than one of the conditions in Theorems 1.2, 1.4 and 1.5.
Proof of Theorem 1.1 $(1)$ and $(2)$ for nilpotent groups.
Let $G$ be a (non-trivial) connected Lie group without any non-trivial compact connected central subgroup, i.e. $G$ belongs to class ${\mathcal{C}^{\prime}}$. Let $T\in{\rm Aut}(G)$.
[…]
As $G$ is simply connected and nilpotent, by Lemma 3.6, ${\rm d\,}T$ acts distally on ${\rm Sub}^{p}_{\mathfrak{g}}$. Then from (1) above, we have that ${\rm d\,}T\in{\mathcal{K}}\{r{\rm Id}\mid r\neq 0\}$ for some compact subgroup ${\mathcal{K}}$ of ${\rm GL}({\mathfrak{g}})$ and ${\rm Id}$ is the identity map on ${\mathfrak{g}}$. Here, ${\rm d\,}T=\psi\,\circ\,r\,{\rm Id}$ for some $r\in{\mathbb{R}}\setminus\{0\}$ and $\psi\in{\mathcal{K}}$. If $r=\pm 1$, then ${\rm d\,}T^{2}\in{\mathcal{K}}$ and $T$ generates a relatively compact subgroup of ${\rm Aut}(G)$. If possible suppose $r\neq\pm 1$. Since ${\rm Aut}(G)$ is almost algebraic, there exists $n\in{\mathbb{N}}$ such that ${\rm d\,}T^{n}$ is contained in a connected almost algebraic subgroup ×${\mathbb{R}}\times\mathbb{T}^{k}$ of ${\rm Aut}(G)^{0}$ (which is identified with its image in ${\rm GL}({\mathfrak{g}})$), where $\mathbb{T}^{k}=(\mathbb{S}^{1})^{k}$ for some $k\in{\mathbb{N}}$. Therefore, $\psi^{n}\,\circ\,r^{n}\,{\rm Id}\in{\mathbb{R}}\times\mathbb{T}^{k}\subset{\rm Aut}(G)^{0}$. Hence both $\psi^{n}$ and $r^{n}\,{\rm Id}\in{\rm Aut}(G)^{0}$. But $r^{n}\,{\rm Id}$ is not a Lie algebra automorphism of ${\mathfrak{g}}$ as $r^{n}\neq 1$; otherwise $r^{n}[X,Y]=r^{n}{\rm Id}[X,Y]=[r^{n}{\rm Id}(X),r^{n}{\rm Id}(Y)]=r^{2n}[X,Y]$, then $[X,Y]=0$ for all $X,Y\in{\mathfrak{g}}$, which leads to a contradiction as $G$ is not abelian. Thus, $r=\pm 1$ and $T$ is contained in a compact subgroup of ${\rm Aut}(G)$. The converse statement is obvious. Thus (2) holds if $G$ is nilpotent. ∎
We will continue with the proof of Theorem 1.1 after proving Theorems 1.2 and 1.3.
@@ 4. Proofs of the main results
Let $X\in{\mathfrak{g}}$, $X\neq 0$ be such that ${\rm d\,}T(X)=X$ and let $A=\{\exp tX\}$, a one-parameter subgroup, it is either closed or its closure $\bar{A}$ is compact. Since $T|_{A}={\rm Id}$, we have that $T|_{\bar{A}}={\rm Id}$ and it follows that $\bar{A}$ normalises $C(T)$ and $\bar{A}\cap C(T)=\{e\}$. Let $H=\overline{A\ltimes C(T)}$. Then $H$ is solvable. As $A$ normalises $C(T)$, so does $H$. Now $H/C(T)$ is connected and abelian as it is the closure of the image of the one-parameter subgroup $A$ in it. Then $H/C(T)$ is either isomorphic to ${\mathbb{R}}$ or it is compact. In the first case it follows that $H=A\ltimes C(T)$. In the second case, since $C(T)$ is a vector group, $H=K\ltimes C(T)$ for some (maximal) compact subgroup $K\subset H$ (cf. [25], Lemma 3.7). If $A$ (resp. $\bar{A}$) is compact, then $H=A\ltimes C(T)$ (resp. $H=\bar{A}\ltimes C(T)$). Here, $H$ is a closed $T$-invariant subgroup of $G$ and $T(h)\in h\,C(T)$ for all $h\in H$.
Step 3: Suppose $H/C(T)$ is isomorphic to ${\mathbb{R}}$. Then so is $A$ and $H=A\ltimes C(T)$. Let $U,V$ be open neighbourhoods of the identity $e$ in $H$ such that $V^{2}\subset U$ and every element of $U$ is contained in a one-parameter subgroup. Let $x=ac$ for some $a\in A\cap V$ and $c\in C(T)\cap V$ with $a\neq e$ and $c\neq e$. Then $x\in V^{2}\subset U$ and $T(x)=xy$ for $y=c^{-1}T(c)\in C(T)$ and $y\neq e$ as $c\in C(T)$ and $c\neq e$. Now $x$ is contained in a one-parameter subgroup (say) $\{x_{t}\}$ in $H$ as $x=x_{1}$. Then $T(x_{t})=x_{t}y_{t}$, $y_{t}\in C(T)$, $t\in{\mathbb{R}}$ and $y_{1}=y\neq e$. Here, $\{x_{t}\}$ is closed as $H$ is simply connected and solvable. As ${\rm Sub}^{p}_{H}$ is compact, we get that $T^{n_{k}}(\{x_{t}\})\to B$ for some $B\in{\rm Sub}^{p}_{H}$ and some unbounded sequence $\{n_{k}\}\subset{\mathbb{N}}$. By Lemma 2.2, $B\neq\{e\}$. Let $b\in B\setminus\{e\}$. Then $T^{n_{k}}(x_{t_{k}})\to b$ for some sequence $\{t_{k}\}\subset{\mathbb{R}}$. Here, $\{x_{t}\}C(T)/C(T)=H/C(T)$ is isomorphic to ${\mathbb{R}}$. Since $x_{t_{k}}C(T)\to b\,C(T)$ in $H/C(T)$, we get that $\{t_{k}\}$ is bounded. Passing to a subsequence if necessary, we may assume that $t_{k}\to t_{0}$. Then $x_{t_{k}}\to x_{t_{0}}$ and $y_{t_{k}}\to y_{t_{0}}$. As all the eigenvalues of $T|_{C(T)}=T_{1}=\phi\circ\alpha\,I$ have absolute value $\alpha$ with $10<\alpha<$ 1, we get that $I-T_{1}$ is invertible in ${\rm GL}(d,{\mathbb{R}})$, where $I$ is the identity matrix in ${\rm GL}(d,{\mathbb{R}})$. Using the addition as the group operation in ${\mathbb{R}}^{d}$, and the fact that $T_{1}^{n}(z)\to 0$ for all $z\in{\mathbb{R}}^{d}$, we get that
@@
Step 4: Now we have that $H/C(T)$ is compact and $H=K\ltimes C(T)$, where $K$ is a maximal compact subgroup of $H$. We show that $T(x)=x$ for all $x\in K$. We know that $T(x)\in x\,C(T)$ for all $x\in K$. Since $K$ is connected and abelian, the set $F$ of roots of unity in $K$ is dense in $K$ and it is enough to show that $T(x)=x$ for all $x\in F$. If possible, suppose there exists $x\in F$ such that $T(x)\neq x$. Let $m\in{\mathbb{N}}$ be the order of $x$, i.e. $x^{m}=e$, and $m\neq 1$. Let $\{x_{t}\}$ be a one-parameter subgroup in $K$ such that $x_{1}=x$. As $x_{1}^{m}=e$, $\{x_{t}\}$ is closed and isomorphic to ${\mathbb{S}}^{1}$ and $T(x_{t})=x_{t}y_{t}$, where $y_{t}\in C(T)$, $t\in{\mathbb{R}}$ and $y_{1}=x^{-1}T(x)\neq e$. There exists an unbounded sequence $\{n_{k}\}\subset{\mathbb{N}}$ such that $T^{n_{k}}(\{x_{t}\})\to B$ for some $B\in{\rm Sub}^{p}_{H}$. Note that $B\neq\{e\}$ (cf. Lemma 2.2). Let $b\in B$ be such that $b\neq e$. Then $T^{n_{k}}(x_{t_{k}})\to b$ for some $\{t_{k}\}\subset{\mathbb{R}}$. Here, $\{x_{t}\}$ is periodic i.e. $x_{mn}=e$ and $x_{m+t}=x_{t}$ for all $n\in{\mathbb{Z}}$ and $t\in{\mathbb{R}}$, where $m$ is the order of $x_{1}=x$. Hence we can choose $\{t_{k}\}$ such that it is bounded. Then $\{x_{t_{k}}\}$, and hence, $\{y_{t_{k}}\}$ is bounded. Now arguing as in Step 3, we get that $T(b)=b$. Therefore, $T(B)=B$, and hence $T^{n_{k}}(T(\{x_{t}\}))\to B$. Since $T$ acts distally on ${\rm Sub}^{p}_{G}$, we get that $T(\{x_{t}\})=\{x_{t}\}$. As $T$ is an automorphism and $\{x_{t}\}$ is isomorphic to ${\mathbb{S}}^{1}$, we get that $T(x_{t})=x_{t}$ or $T(x_{t})=x_{t}^{-1}=x_{-t}$. In either case, we get that $y_{1}\in K\cap C(T)=\{e\}$. This leads to a contradiction. Therefore, $T(x)=x$, for all $x\in F$, and hence for all $x\in K$.